use multiplication with the addition-or-subtraction method to solve the system:
5x + 2y = 14
3x - 5y =27

Answer by mananth(16949)

5x+2y=14 .............1
3x-5y=27 .............2
Eliminate y
multiply (1)by 5
Multiply (2) by 2
25x+10y= 70
6x-10y= 54
Add the two equations
31x=124
/31
x=4.00
plug value of x in (1)
5x+2y=14
20+2y=14
2y=14-20
2y=-6
y=-3.00